[DMe] Define group. Show that the set P3 of all permutations on three symbols 1,2,3 is a finite non-abelian group of order six with respect to permutation multiplication as composition.


1
Expert's answer
2022-02-17T10:06:17-0500

Solution:

A group is a set G, combined with an operation *, such that:

  1. The group contains an identity
  2. The group contains inverses
  3. The operation is associative
  4. The group is closed under the operation.


Here "P_3=\{I,(1\ \ 2), (2 \ \ 3), (3 \ \ 1), (1 \ \ 2\ \ 3), (1 \ \ 3\ \ 2)\}"  where "\mathrm{I}"  is the identity permutation.

There are six elements in the set P3 .

Let "\quad f_{1}=I, f_{2}=\left(\begin{array}{ll}1 & 2\end{array}\right), f_{3}=\left(\begin{array}{ll}2 & 3\end{array}\right), f_{4}=\left(\begin{array}{ll}3 & 1\end{array}\right), f_{5}=\left(\begin{array}{lll}1 & 2 & 3\end{array}\right)\ and\ f_{6}=\left(\begin{array}{lll}1 & 3 & 2\end{array}\right)"

Let us prepare a composition table for P3 .



For completing the entries in the above table we have actually multiplied the permutations. Thus

"\begin{aligned}

&f_{2} f_{3}=\left(\begin{array}{lll}

1 & 2

\end{array}\right)\left(\begin{array}{ll}

2 & 3

\end{array}\right)=\left(\begin{array}{lll}

1 & 3 & 2

\end{array}\right)=f_{6} \\

&f_{2} f_{4}=\left(\begin{array}{lll}

1 & 2

\end{array}\right)\left(\begin{array}{ll}

3 & 1

\end{array}\right)=\left(\begin{array}{lll}

1 & 2 & 3

\end{array}\right)=f_{5} \\

&f_{3} f_{3}=\left(\begin{array}{lll}

2 & 3

\end{array}\right)\left(\begin{array}{ll}

2 & 3

\end{array}\right)=f_{1}, \text { the identity permutation. } \\

&f_{5} f_{5}=\left(\begin{array}{llll}

1 & 2 & 3

\end{array}\right)\left(\begin{array}{lll}

1 & 2 & 3

\end{array}\right)=\left(\begin{array}{lll}

1 & 3 & 2

\end{array}\right)=f_{6} \text { etc. }

\end{aligned}"

Now we observe that

"\left(G_{1}\right)"  All entries in the table belong to "P_{3}"  and therefore "P_{3}"  is closed with respect to product of permutations.

"\left(G_{2}\right)" Product of permutations is an associative operation.

 "\left(G_{3}\right)" The identity permutation f1 is the identity of multiplication.

 "\left(G_{4}\right)"  Every element of P3 possesses inverse in P3 because

inverse of f1 is f1 ; inverse of f2 is f2 ; inverse of f3 is f3

inverse of f4 is f4 ; inverse of f5 is f5 ; and inverse of f6 is f6 .

"\left(G_{5}\right)"  The product of permutations in P3 is not commutative as

"f_{2} f_{3}=f_{6}\ and\ f_{3} f_{2}=f_{5}\ and\ f_{2} f_{3} \neq f_{3} f_{2}" .

Therefore P3  is a finite non-abelian group of order six with respect to permutation multiplication.